Status Report

NOAA GOES SATOPS Morning Report 6 Feb 2001

By SpaceRef Editor
February 6, 2001
Filed under ,

SATOPS MORNING REPORT February 6, 2001

 

POES

* Frame Synchronizer testing continues.

* NOAA-16 recorder (DTR5B) removed from spacecraft memory loads until
further notice.  Recorder continuing to have data missing when played
back.  Anomaly resolution continues.

NOAA-16 rev 1937 / W at 1824z on February 5: Uplinked commands to power-on
VTX2 and select APT channels 2 & 4.

NOAA-11 rev 63784 / F at 2220z on February 5:  Only 90% of S1A
was received from the spacecraft and CEMSCS reported that their equipment
received it with a NOAA-F  header.  CEMSCS monitor in SOCC showed
the correct header, so a replay was not requested.

NOAA-12 rev 50556 / W at 1006z on February 5: G4B playback short with
87.2% recovered.  Recovery attempt will be made on rev 50559.

 

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

 

GOES

No change in the status of the GOES spacecraft.

GOES-8 Rapid Scan Operations requested by NWS/Pittsburgh for winter
storm in New England and approved for 1626z February 5 to 0126z February
6.

At 0835z on February 6: GOES-8 image frame break due to loss of signal
synch.

At 0302z on February 6: GOES-10 image frame break due to loss of signal
synch.

 

 

 

Questions or comments to msimpson@nesdis.noaa.gov

SpaceRef staff editor.